OpenStudy (pottersheep):

PLEASE HELP. Is this correct? cos(90 +Q ) = -sinQ PLEASE HELP. Is this correct? cos(90 +Q ) = -sinQ @Mathematics

OpenStudy (anonymous):

Yes..

OpenStudy (jamesj):

Yes, it is. And you can prove it using the identity: cos(x + y) = cos x . cos y - sin x . sin y With x = 90 and y = Q, we have cos(90 + Q) = cos 90 . cos Q - sin 90 . sin Q = 0 . cos Q - 1 . sin Q = - sin Q

OpenStudy (pottersheep):

Thanks!